OpenStudy (anonymous):

6.find the 9th term of the sequence described by :A(n)=7+(n-1)(3). A)42 B)9 C)31 D)27 7.What is the function for the arithmetic sequence 2,7,12,17,22. A) A(n)=5+(n-1)(2) B) A(n)=2+(n-1)(6) C) A(n)=5(6) D A(n)=2+(n-1)(5)

OpenStudy (mathmath333):

\(\large\tt \color{black}{A(n)=7+(n-1)(3)}\) \(\large\tt \color{black}{put~~n=9}\)

OpenStudy (mathmath333):

\(\large\tt \color{black}{2,7,12,17,22}\) \(\large\tt \color{black}{a=2,d=7-2=12-7=17-12=5,}\) \(\large\tt \color{black}{A(n)=a+(n-1)(d)}\)

OpenStudy (anonymous):

so 6. is B? and 7. is D?

OpenStudy (mathmath333):

7. is correct

OpenStudy (mathmath333):

u should recalculate 6

OpenStudy (anonymous):

oh ok so its C 31 if you plug in 9 in the for the n so its 7+(9-1)(3)=31 right?

OpenStudy (mathmath333):

yes of cos

OpenStudy (anonymous):

ok thank you so much for all of you help

OpenStudy (mathmath333):

\(\Huge\color{blue}{\ddot\smile}\)